### v1

# Can Input Attributions Interpret the Inductive Reasoning Process Elicited in In-Context Learning?

Abstract
Elucidating the rationale behind neural models’ outputs has been challenging in the machine learning field, which is indeed applicable in this age of large language models (LLMs) and in-context learning (ICL). When it comes to estimating input attributions (IA), ICL poses a new issue of interpreting which example in the prompt, consisting of a set of examples, contributed to identifying the task/rule to be solved. To this end, in this paper, we introduce synthetic diagnostic tasks inspired by the poverty of the stimulus design in inductive reasoning; here, most in-context examples are ambiguous w.r.t. their underlying rule, and one critical example disambiguates the task demonstrated. The question is whether conventional IA methods can identify such an example in interpreting the inductive reasoning process in ICL. Our experiments provide several practical findings; for example, a certain simple IA method works the best, and the larger the model, the generally harder it is to interpret the ICL with gradient-based IA methods. [^1] We will make our data and scripts public upon the publication of this paper.
[…]
github   https://github.com/muyo8692/input-attribution-icl
Can Input Attributions Interpret the Inductive Reasoning Process
Elicited in In-Context Learning?
Mengyu Ye1     Tatsuki Kuribayashi2     Goro Kobayashi1,3     Jun Suzuki1 1Tohoku University     2MBZUAI     3RIKEN {ye.mengyu.s1, goro.koba}@dc.tohoku.ac.jp tatsuki.kuribayashi@mbzuai.ac.ae     jun.suzuki@tohoku.ac.jp
Figure 1: Overview of our experimental setup. The majority of in-context examples (gray) are ambiguous, supporting either Rule-A of adding two tokens or Rule-B of doubling tokens. A single disambiguating example (blue) reveals that Rule-A is correct. We investigate whether Input Attribution (IA) methods can track such an inductive reasoning process.
1 Introduction
In past years, input attribution (IA) methods, e.g., gradient norm Simonyan et al. (2014); Li et al. (2016a), have typically been employed in the natural language processing (NLP) field to interpret input–output associations exploited by neural NLP models Vinyals and Le (2015); Li et al. (2016b). Such IA approaches are now somewhat outdated because of the rise of large language models (LLMs) and mechanistic interpretability (MI) communities Olah et al. (2020); Bereska and Gavves (2024), where the general aim is to understand the internal mechanism of LLMs and control their behaviors by intervening in their internal representations or weight parameters rather than elucidating their input-output associations. Despite the enriched scope of modern interpretability research, such rapid progress has missed some intriguing questions bridging the IA and MI eras: in particular, do conventional IA methods still empirically work in the modern NLP setting, specifically in the context of LLMs and in-context learning (ICL)?
In this paper, we perform a systematic, controlled evaluation of IA methods in interpreting the inductive reasoning process of LLMs in ICL Brown et al. (2020)Can IA methods interpret the example in a few-shot prompt that most influenced LLMs’ inductive learning of the task? This question is worth investigating for several reasons. First, input attribution would still serve as a necessary and sufficient explanation in typical practical cases. Some users might simply seek which part of the context is heavily referred to by an LLM system; here, its detailed internal processes identified by MI methods will be rather redundant.
Second, the modern NLP setting, specifically ICL, differs from the conventional tasks where IA methods have been tested. Simply put, in the conventional settings, the task objective is given (e.g., sentiment analysis) in advance, during training, and the IA is employed to just identify the association between input $X$ (i.e., tokens) and output $y$. In contrast, in ICL, the model should (i) first identify the task solely from limited examples demonstrated and then (ii) answer a new question; the former step particularly introduces a new aspect to be interpreted by IA methods — which example, the pair of $(X_{k},y_{k})$ in input, contributed to identifying the task to be solved for a model. This is a type of instance-based interpretation of neural models Wachter et al. (2017); Charpiat et al. (2019); Hanawa et al. (2021), and such an in-context, inductive learning process of LLMs attracts interest from a scientific perspective as well Treutlein et al. (2024); Mueller et al. (2024).
[…]
Our experiments reveal several findings:
  • Certain simpler IA methods (input erasure, gradient norm) frequently outperform other IA methods (e.g., Integrated Gradient) and other interpretability approaches such as observing attention weights, suggesting the particular (dis)advantages of existing IA methods under the ICL×$\times$×LLM settings.
  • Gradient-based and other IA methods did not work stably across different tasks and models, posing their general limitations in interpreting the inductive reasoning part of ICL.
  • Different approaches exhibited different properties against the scaling of the experimental setting; for example, IA methods perform better in many-shot scenarios, whereas a particular baseline (self-answer) tends to outperform IA in larger models.
2 Preliminary
2.1 Input attribution (IA) methods
Input attribution (IA) methods Denil et al. (2014); Li et al. (2016a); Poerner et al. (2018); Arras et al. (2019); Yin and Neubig (2022) are one of the most popular and traditional techniques for interpreting and explaining the predictions of machine learning models. Specifically, IA methods determine how much each input feature contributes to a particular prediction; that is, given input tokens $X\coloneqq[x_{1},\ldots,x_{n}]$ and output $y$, the IA methods yield the strength of contribution $S(x_{i})$ of each input $x_{i}$ to the output $y$. Note that the input in ICL consists of several in-context examples (§ 2.2); here, we represent all the tokens in an input sequence as $X$, and the answer to the target question is denoted as $y$. We examine the following four representative IA methods in the ICL×$\times$×LLM context:
Input erasure (IE)
IE Li et al. (2016c) measures how erasing a certain token $x_{i}$ from the input prompt makes the target token $y_{t}$ less likely to be generated:
$S_{\mathrm{IE}}(x_{i},y_{t};X)$ $=q(y_{t}|\bm{X})-q(y_{t}|\bm{X}_{\neg i})\text{,}$ (1)
where $\bm{x}_{i}\in\mathbb{R}^{d}$ denotes the $d$-dim input token embedding for the $i$-th token in an input, and $\bm{X}\coloneqq[\bm{x}_{1},\ldots,\bm{x}_{n}]$ denotes the sequence of input token embeddings. $\bm{X}_{\neg i}\coloneqq[\bm{x}_{1},\ldots,\bm{x}_{i-1},\bm{x}_{i+1},\ldots,\bm{x}_{n}]$ denotes the sequence of input token embeddings without $\bm{x}_{i}$. We emulate this partial input $\bm{X}_{\neg i}$ by introducing an attention mask to zero-out the attention to $\bm{x}_{i}$ in every layer (thus, the original position information holds). $q(y|\bm{X})$ represents the model’s prediction probability for the token $y_{t}$ given input $\bm{X}$.
Gradient norm (GN)
GN Simonyan et al. (2014); Li et al. (2016a) calculates the attribution (contribution) score for each input token $x_{i}$ by computing the L1 norm of its gradient of the target token $y_{t}$:
$S_{\mathrm{GN}}(\bm{x}_{i},y_{t};\bm{X})={\lVert g(\bm{x}_{i},y_{t};\bm{X})\rVert}_{\mathrm{L1}}$ (2)
$g(\bm{x}_{i},y_{t};\bm{X})=\nabla_{\bm{x}_{i}}q(y_{t}|\bm{X})\text{.}$ (3)
where $g(\bm{x}_{i},y_{t};\bm{X})\in\mathbb{R}^{d}$ denotes the gradient computed for $\bm{x}_{i}$, given an output $y_{t}$ and input embedding sequence $\bm{X}$.
Input×$\times$×gradient (I×$\times$×G)
I×$\times$×G Shrikumar et al. (2017); Denil et al. (2014) takes the dot product of the gradient with the input token embedding $\bm{x}_{i}$:
$S_{\mathrm{I\times G}}(\bm{x}_{i},y_{t};\bm{X})=g(\bm{x}_{i},y_{t};\bm{X})\cdot\bm{x}_{i}\text{.}$ (4)
Integrated gradients (IG)
IG Sundararajan et al. (2017) computes attribution scores by accumulating gradients along a straight path from a baseline input $\bm{X}^{\prime}$ to the actual input $\bm{X}$:
$S_{\mathrm{IG}}(\bm{x}_{i},y_{t};\bm{X})=$
[…]
where $\bm{X}^{\prime}\coloneqq[\bm{x}^{\prime}_{1},\ldots,\bm{x}^{\prime}_{n}]$ denotes the sequence of baseline embeddings222 We followed the common practice and employed a sequence of zero vectors as the baseline input. We used an interpretability library captum Kokhlikyan et al. (2020) to calculate the IG score and keep all parameters as default., and $\alpha$ denotes the interpolation coefficient. In practice, the integral is approximated using numerical integration with a finite number of steps.
Contrastive explanations
For the IE, GN, and I×$\times$×G methods, we also adopt the contrastive explanation setting Yin and Neubig (2022). IA methods in this setting measure how much an input token $x_{i}$ influences the model to increase the probability of target token $y_{t}$ while decreasing that of foil token $y_{f}$. A foil token can be defined as an output with an alternative, incorrect generalization (§ 3). Contrastive versions of IE, GN, and I×$\times$×G are defined as follows:
$\!\!S_{\mathrm{IE}}^{*}(x_{i},y_{t},y_{f};\bm{X})$
$\!\!\qquad=S_{\mathrm{IE}}(x_{i},y_{t};\bm{X})-S_{\mathrm{IE}}(x_{i},y_{f};\bm{X})$ (6)
$\!\!S_{\mathrm{GN}}^{*}(\bm{x}_{i},y_{t},y_{f};\bm{X})=\lVert g^{*}(\bm{x}_{i},y_{t},y_{f};\bm{X})\rVert_{\mathrm{L1}}$ (7)
$\!\!S_{\mathrm{I\times G}}^{*}(\bm{x}_{i},y_{t},y_{f};\bm{X})\!=\!g^{*}(\bm{x}_{i},y_{t},y_{f};\bm{X})\cdot\bm{x}_{i}$ (8)
$\!\!g^{*}(\bm{x}_{i},y_{t},y_{f};\bm{X})=\nabla_{\bm{x}_{i}}\bigl{(}q(y_{t}|\bm{X})-q(y_{f}|\bm{X})\bigr{)}$ (9)
We report the results with contrastive explanation in the main part of this paper § 5.
2.2 Interpreting in-context learning (ICL)
We focus on the ICL setting Brown et al. (2020), which has typically been adopted in modern LLM-based reasoning. An input prompt in ICL setting consists of few-shot examples $E$ and a target question. $E$ is composed of $n$ examples [$e_{1},\cdots,e_{n}$], each of which contains an input-output pair $e_{i}=(X_{i},f(X_{i}))$, given a function $f$ associated with the task. Let $X_{n+1}$ represent the target question $q$ that the model must answer. The ICL setting is formed as follows:
$\underbrace{\overbrace{\overbrace{X_{1},f(X_{1})}^{\textit{example}~{}e_{1}},\ldots,\overbrace{X_{n},f(X_{n})}^{\textit{example}~{}e_{n}}}^{\textit{few-shot examples E}},\overbrace{X_{n+1}}^{q}}_{\textit{prompt}},\underbrace{f(X_{n+1})}_{\textit{completion}}$
Here, a model is expected to (i) first induce the underlying function (rule) $f$ from examples $E$ and then generate the final output $f(X_{n+1})$.
Eureka example
[…]
Instance-level attribution
Notably, we consider the use of IA methods to identify a particular example $e^{*}\in E$ in input, instead of a particular token. To compute an IA score for an example $S(e_{i})$, we sum up the IA scores for its constituent tokens: $S(e_{i})=\sum_{x_{j}\in(X_{i},y_{i})=e_{i}}S(x_{j})$. [^3] An exception applies in the IE method; the attribution score for an example $e_{i}$ is simply computed by erasing the corresponding $X_{i}$ and $f(X_{i})$ from the input sequence. Our interest is which example obtains the highest IA score, i.e., $\operatorname*{arg\,max}_{e_{i}\in E}S(e_{i})$.
Task Prompt example/template Answer Potential rules
Associative-Recall
a $\mapsto$ 6
g $\mapsto$ 3
w $\mapsto$ 5
g $\mapsto$
3
Key–value pairs are in the prompt.
The task is to output a value associated
with a given key.
Linear-or-Distinct
a a b a $\mapsto$ b
g g j g $\mapsto$ j
k i k k $\mapsto$ k / i
o o o p $\mapsto$
o / p
A. Generate the n-th token
     (3rd token in this example)
B. Generate the distinctive token
Add-or-Multiply
aa $\mapsto$ aaaa
hh $\mapsto$ hhhh
vvv $\mapsto$ vvvvv / vvvvvv
i $\mapsto$
iii / ii
A. Multiply the numder of tokens by $n$
($n=2$ in this example)
B. Add $m$ tokens
($m=2$ in this example)
Verb-Object
like [CITY] $\mapsto$ True
love [ANIMAL] $\mapsto$ False
like [ANIMAL] $\mapsto$ True / False
love [CITY] $\mapsto$
False / True
A. If “like” exists, then True
B. If [CITY] exists, then True
Tense-Article
The [NOUN] [VERB]-ing $\mapsto$ True
A [NOUN] [VERB]-past $\mapsto$ False
A [NOUN] [VERB]-ing $\mapsto$ True / False
The [NOUN] [VERB]-past $\mapsto$
False / True
A. If the verb is in ing form, then True
B. If the first token is “the", then True
Pos-Title
The [NOUN] Was [ADJ] $\mapsto$ True
The [noun] was [noun] $\mapsto$ False
The [noun] was [adj] $\mapsto$ True / False
The [NOUN] Was [NOUN] $\mapsto$
False / True
A. If adjective exsist, then True
B. If the sentence is in title case, then True
Table 1: General format of our inductive reasoning tasks. As a baseline setting, we also set Associative-Recall setting to just memorize key-value mappings. The remaining tasks span from somewhat superficial features to linguistic ones. The disambiguating example (the third one in these examples) determines the correct rule and answer (blue or orange) for the final question from two plausible generalizations shown in the “Potential rules” column.
3 Poverty of the stimulus scenarios as testbeds
We evaluate how well a particular IA method can identify the in-context example $e^{*}$ that most contributes to identifying the task—specifically, whether the IA method assigns the highest attribution score to such an example. However, in real-world tasks, it is generally unclear which in-context example is the most influential in disambiguating the task. Moreovdr, in some cases, the task may be solved without relying on any of the examples (e.g., solved by leveraging prior knowledge). Therefore, these are not suitable as a benchmark to evaluate the interpretability method, and we decided to design a synthetic and controlled task setting.
[…]
Linear-or-Distinct (LD)
The few-shot examples are ambiguous as to Rule-A: selecting a character in a particular linear position in an input $X_{i}$; or Rule-B: selecting a character that differs from the others in an input $X_{i}$.
Add-or-Multiply (AM)
The ambiguity of this task is Rule-A: multiply the numbers of tokens in the input $X_{i}$; or Rule-B: add a certain number of tokens to input $X_{i}$.
Verb-Obeject (VO)
This task requires distinguishing whether the controlling condition is the type of verb (Rule A) or the category of the object noun (Rule B). There are two verbs (“like” and “love”) and two categories of the object (city or animal).
Tense-Article (TA)
The potential rules are Rule-A: whether the main verb in the input $X_{i}$ is in ing-form or not; or Rule-B: whether the first token of input $X_{i}$ is “The” or not.
Pos-Title (PT)
This task involves two rules: Rule-A: whether there is an adjective in $X_{i}$; or Rule-B: whether $X_{i}$ is presented in title case.
Figure 2: Attribution accuracies for each task/model in the 10-shot setting (thus, the chance rate is 10%; red dotted line). The edit distance and attention baselines are indicated by a black dotted line and gray bar. respectively.
In addition to them, we adopted a simple task of associative recall (AR), which is typically employed in studying ICL, where the model is supposed to simply memorize the key:value mapping rules demonstrated in the prompt and apply them to the target question. Linguists may be more interested in the task of, for example, syntactic transformation to an interrogative sentence McCoy et al. (2020) based on the original poverty of the stimulus argument in the language domain Chomsky (1980). However, such a realistic task interferes with the models’ meta-linguistic knowledge; thus, we adopted artificial (and somewhat simpler) ones.
Foil token
A contrastive explanation needs a foil token corresponding to an explicit negative label (§ 2.1). We use the token/answer corresponding to an alternative rule (conflicting the disambiguating example) as the foil token.
4 Experimental setup
4.1 Overview
Few-shot settings
We conduct experiments with different numbers of few-shot examples; specifically, we examined 10-shot, 50-shot, and 100-shot settings to test the robustness of IA methods toward somewhat longer demonstrations.
Data
For each synthetic task, we create 200 different questions with different sets of few-shot examples and a target question. In the LD, AM, VO, TA, and PT tasks, the correct rule is selected out of the two candidates (rules A or B shown in Table 1) in a 1:1 ratio. The position of the most influential (i.e., disambiguating) example $e^{*}$ is controlled to be uniformly distributed by chance. We test IA methods using only the questions that models answered correctly.
Metrics
We report the ratio of questions where the most influential example $e^{*}$ is picked by the highest IA score, i.e., how frequently $\operatorname*{arg\,max}_{e_{i}\in E}S(e_{i})$ is equal to $e^{*}$. Henceforth, we call this score attribution accuracy.
Models
We evaluate five LLMs: Llama-2-7B, Llama-2-13B Touvron et al. (2023), Gemma-2-2B, Gemma-2-9B, and Gemma-2-27B Rivière et al. (2024). As a prerequisite of our experiments, the models should be able to learn the task, i.e., be sufficiently sensitive to the disambiguating example $e^{*}$ and correctly use this to determine the rule. To ensure this ability, we fine-tune these models on each task (see Appendix A), but the conclusions overall did not alter before and after fine-tuning (see Appendix B).
4.2 Baselines
Along with the IA methods introduced in § 2.1, we evaluate four baseline methods.
Edit distance
This method identifies $e^{*}$ simply using edit distance between the target example $X_{n+1}\oplus y_{n+1}$ and each example $X_{i}\oplus y_{i}$, where $\oplus$ is a string concatenation. Example with the minimum edit distance, thus the most similar example to the target question, is selected as an explanation. The weak performance of this problem probes that our experimental setting is so challenging that just relying on surface features does not resolve it.
Attention weights
This method leverages attention weights, computed as the sum of attention weights across all tokens in input $X$. While attention weights are generally considered unreliable for model interpretation, we include this baseline to compare whether IA methods achieve superior performance.
Self-answer
We also examine a somewhat naive approach of directly asking the models to generate their rationale. Specifically, we have models generate the most informative example in a prompt (Appendix C) in deriving their answer.
Chance rate
We also report the chance rate of attribution accuracy when randomly selecting one example from a prompt.
5 Experiments
Figure 2 shows the results in 10-shot settings. [^4] Additional results using fine-tuned models and non-contrastive IA methods are provided in Appendix D.
5.1 Main results
Input erasure works well
The IE method yielded the best attribution accuracy in identifying the most informative example in most cases. However, it is computationally costly since it requires multiple forward passes with different inputs ($O(n^{3})$ compared to the standard $O(n^{2})$ complexity). That is, if the computational costs are not the problem, simply applying the input erasure is the most reliable way to estimate the importance of each example in ICL. At the same time, despite our controlled task settings and nearly 100% accuracy of LLMs in yielding the correct answer, the IE method does not yield the perfect attribution accuracy in some tasks; this limitation is discussed in § 6. Notably, baselines, such as edit distance of the examples, did not yield a good attribution accuracy overall, which proves that our test suites are challenging.
Improved versions of gradient norm do not show advantage
Among the gradient-based methods, simple gradient norm tends to work the best in most (i.e., LD, AM, VO, SP, ST) tasks. In other words, whereas I $\times$ G and IG are proposed as the improved version of simple gradient norm method, there were no substantial advantages of these methods in our settings. In particular, IG consistently yielded the lowest attribution accuracy across all six tasks among the gradient-based methods, suggesting its limitations in ICL scenarios. The plausible reason behind this inferiority is discussed in § 6.
5.2 Scaling properties
LLMs have progressively been scaled up toward model parameter size and its context length in the NLP community. We analyze how such a scaling affects the LLMs’ interpretability, particularly in our setting.
Interpretability vs. model size
We first investigate the relationship between attribution accuracy and model size — is it more difficult to interpret larger LLMs? We observe somewhat intriguing patterns for this question (Figure 5); gradient-based methods tend to work worse in larger LLMs, and in contrast, the self-answer baseline works better in larger LLMs. That is, the (empirically) appropriate approach to interpret the LLMs may differ in their model scale.
Figure 3: Attribution accuracy across six tasks for different Gemma-2 model sizes. Results demonstrate that gradient-based attribution methods show declining performance in larger LLMs, while the self-answer exhibits improved performance with increasing model size.
Figure 4: Attribution accuracy for interpreting Gemma-2-2B models across all six tasks. Gradient-based methods is relatively robust to the number of few-shot examples, while there is a consistent, large drop in attribution accuracy in SA. Note that both x-axis and y-axis are in log scale, and we did not conduct experiments on 100-shot due to computation resource limitations.
Interpretability vs. number of examples
Next, given the trend of long-context LLMs, we examine the relationship between attribution accuracy and the number of few-shot examples. Figure 4 shows the attribution accuracy for Gemma-2-2B across all six tasks. This demonstrates that gradient-based methods maintain accuracy or rather improve against the longer context, in contrast to the decreasing chance rate. [^5] For IG in this setting, due to computation resource limitations, we did not conduct experiments. This suggests the robustness of IA methods in long-context scenarios, highlighting their potential for interpreting inputs with extensive contextual information. Notably, the quality of self-answer consistently degraded as the number of in-context examples increased, while the positive scaling effect observed in the previous analysis of model size. That is, the gradient-based methods and self-answer approach exhibit an insightful trade-off between different scaling properties.
Task Acc. (%) IE Worked?
LD (a) $98.0$ No
LD (b) $0.5$
AM (a) $34.5$ Yes
AM (b) $65.5$
VO (a) $100.0$ No
VO (b) $0.0$
SP (a) $98.0$ No
SP (b) $2.0$
ST (a) $68.5$ No
ST (b) $31.5$
Table 2: Task accuracy (not attribution accuracy) of Gemma-2-2B (excluding AR) when the disambiguating example is not included, separated by the correct rule. The accuracy drastically differs when the correct rule is different; thus, the models adopt a particular default rule with their inductive biases against fully ambiguous demonstrations, even in our controlled settings.
6 Discussion
This section discusses the potential reasons for the unexpected results presented in § 5, highlighting the challenging issues in interpreting ICL.
Why did IE fail to ahieve 100% attribution accuracy?
Our tasks can not be answered without disambiguating examples. Thus, it is somewhat unintuitive to see the non-100% attribution accuracy of the IE method (again, the LLMs understood the task as they achieved 100% accuracy in the tasks) — what happens here? To obtain a hint to clarify IE’s potential limitations, we analyze model behaviors when the disambiguating example is excluded. Interestingly, LLMs adopted a specific generalization (rule) in each task when there was no disambiguating example (Table 2); in other words, they sometimes exhibited strong inductive biases in our tasks. That is, when the correct rule is equal to their preferred rule by their inductive bias, they can answer the task correctly even without disambiguating examples, and the IE method does not compute a proper attribution score. It is now common to see LLMs have particular inductive biases (not a tabula rasaWarstadt et al. (2020); Kharitonov and Chaabouni (2021). Catching such generalization bias with IA methods represents an inherent challenge, highlighting their potential limitations in interpreting ICL.
Why was I×$\times$×G worse than gradient norm?
One advantage of the I×$\times$×G method, compared to the simple gradient norm, is the consideration of the norm of the input embedding Shrikumar et al. (2017); Denil et al. (2014). Since a large vector tends to have a large dot product with another vector, the norm of the input token embedding (vector) is expected to affect the IA score of I×$\times$×G. Then, no improvement of I×$\times$×G over the simple gradient norm suggests that, at least in our settings, the norm of the embeddings was not informative to estimate the input attribution. The norm of the embedding largely has decontextualized information about the word, such as frequency, and it may make sense that such information is not helpful to interpreting our controlled, artificial ICL tasks consisting of alphabet characters, numbers, or random words.
Why was IG worse than gradient norm?
IG is a path-based approach; the gradient is accumulated from a baseline vector (typically zero vector) to the targeted input representation (in our case, the sequence of input embeddings representing few-shot examples). This approach is somewhat intuitive when considering an attribution for a particular word or sentence; for example, suppose one computes an attribution to the word “excellent” in an input for a particular task-specific model, IG may trace the path from zero to the “excellent” vector, which will be in a kind of the goodness direction, involving the points corresponding to, e.g., “okay” “decent,” “good,” “excellent” Sanyal and Ren (2021). Then, one critical question is — what does this path mean in the prompt/task space? Different prompt representations will no longer correspond to the same task; thus, the attribution of a particular token in the middle of such a path in the prompt space may no longer be an attribution under a targeted task. This can be one concern toward the ineffectiveness of IG in our settings.
7 Related Work
IA methods
Several lines of research are conducted to interpret neural language models. NLP researchers have adapted IA methods, which was originally applied to vision models Simonyan et al. (2014); Springenberg et al. (2015); Zintgraf et al. (2017), to perform a post-hoc interpretation of input-output associations exploited by language models Karpathy et al. (2015); Li et al. (2016a); Arras et al. (2016); Lei et al. (2016); Alvarez-Melis and Jaakkola (2017), and its improved versions have also been developed Denil et al. (2014); Sundararajan et al. (2017); Murdoch et al. (2018); Sinha et al. (2021); Ding and Koehn (2021); Bastings et al. (2022); Yin and Neubig (2022); Ferrando et al. (2023). In line with these studies, we provide a new perspective to evaluate these IA methods in ICL. Note that, as an orthogonal attempt, some research estimates the saliency scores to directly prompt models to generate such explanations Rajani et al. (2019); Liu et al. (2019); Wu and Mooney (2019); Narang et al. (2020); Marasovic et al. (2022). This method is indeed examined as one baseline in our study.
Instance-based explanation
Instance-based explanation seeks for the explanation in training data rather than the immediate input during inference as IA methods Wachter et al. (2017); Charpiat et al. (2019); Hanawa et al. (2021). These two paradigms of instance-based and IA-based explanations have been studied somewhat separately since the information source to seek the explanation is clearly different. On the other hand, in ICL, the training examples are now in the input during inference that can be analyzed by IA methods. In this sense, our investigation can seen as a new exploration of instance-based explanation with the help of IA methods.
Mechanistic Interpretability
With the rise of large language models, such as GPT-3 Brown et al. (2020), the mechanistic interpretability community has shifted its focus from vision models to language models. Within which, the promising results using sparse autoencoders (SAEs) Bricken et al. (2023); Templeton et al. (2024) have inspired a flurry of follow-up work Gao et al. (2024); Lieberum et al. (2024); Rajamanoharan et al. (2024a, b); Karvonen et al. (2024); Braun et al. (2024); Kissane et al. (2024); Makelov (2024). Such a scope of SAE, interpreting the model internals, is orthogonal to our direction of estimating the importance of input examples.
8 Conclusions
We have pointed out and tackled the problem of interpreting the inductive reasoning process in ICL as a missing but reasonable milestone to be explored in LLM interpretability research. Our revisit to the IA methods in interpreting this ICL process has clarified their limitations from a new angle as well as provided fruitful insights and discussions on their practical usage in modern NLP. These findings have highlighted some issues in the community; in particular, even the fundamental task of mapping input and output has not been accomplished, and there is room to sophisticate previously developed interpretability tools to be suitable for LLMs.
Limitations
Our study has several limitations in scope. First, we focused primarily on popular gradient-based IA methods, leaving other approaches such as perturbation-based methods like LIME Ribeiro et al. (2016) and SHAP Lundberg and Lee (2017) for future work.
Regarding model selection, we concentrated on widely-used open-weight LLMs. Since applying IA methods requires gradient computation through backward propagation, computational constraints limited our ability to evaluate all available models, particularly large ones such as Llama-2-70B Touvron et al. (2023) and Llama-3.1-405B Dubey et al. (2024), and limited us to evaluate IG in § 5.2.
Our experimental design used synthetic tasks to better identify influential examples in the few-shot setting. While this approach allowed for controlled experimentation, both the number and format of tasks were limited. Future work could explore more realistic tasks with greater variations.
We focused exclusively on pre-training models, excluding post-training models from our analysis. This choice was motivated by our interest in basic few-shot learning, which is more commonly used with pre-trained models. Although post-training models might demonstrate higher accuracy on our tasks and the self-answer setting due to their potentially superior capabilities Rivière et al. (2024), our primary focus was on evaluating IA methods rather than model performance.
The optimization of the self-answer setting was not explored in depth, as our main interest lay in examining whether larger models showed improved performance in this setting rather than enhancing the setting itself.
Finally, while our ambiguous tasks were designed with two potential functions in mind, we acknowledge that models might interpret these tasks differently than intended. However, since our focus was on whether models could recognize these as ambiguous tasks with two possible answers and use specific examples to determine the appropriate response, we believe this limitation does not significantly impact our findings.
Ethical Statements
This work advances our understanding of input attribution (IA) methods in the context of large language models’ (LLMs) in-context learning (ICL). Our findings contribute to the broader goal of developing more interpretable and safer AI systems by providing practical insights into the strengths and weaknesses of IA methods as tools for interpreting LLMs.
This study exclusively uses synthetic data generated through computational methods. No real user data, human annotations, or personally identifiable information was collected or used in our experiments. Our synthetic dataset generation process does not involve any human subjects, crowd workers, or demographic information.
Acknowledgements
We would like to express our gratitude to the members of the Tohoku NLP Group for their insightful comments. And special thanks to Sho Yokoi for his valuable suggestions on how to improve clarity in several aspects. This work was supported by the JSPS KAKENHI Grant Number JP19H04162, JP22J21492; JST Moonshot R&D Grant Number JPMJMS2011 (fundamental research); JST BOOST, Japan Grant Number JPMJBS2421.
Appendix A Finetuning Details
Fine-tuning data consisted of equal numbers of 10-shot, 50-shot, and 100-shot examples.
For each task, we create the train set for fine-tuning use tokens with no overlaping with the test set (the dataset we used for our main experments).
We fine-tuned models on each task seperately, this allowes them to focus solely on learning the underlying rules of that task, without needing to identify which task is being presented during testing
This results in six fine-tuned models per LLM, one for each task, with the exceprtion of Gemma-2-27B model, which we didn’t fine-tune Gemma-2-27B on the Associative-Recall task as the original model performs good enough on this simple task.
A.1 Finetuing parameters
We use a fixed LoRA configuration with rank $r=32$ and scaling factor $\alpha=64$, applying a dropout rate of $0.05$ across all linear modules. The LoRA adaptation includes bias terms in the training.
For optimization, we perform a learning rate sweep using a cosine scheduler with a $5\%$ warm-up period relative to the total training steps. The optimal learning rate typically falls in the order of $1\times 10^{-5}$ when the loss reaches its minimum.
In our experiments, the Llama-2 models achieve zero loss, which is expected in such synthetic setting. The Gemma-2 models, however, converge to final loss values of approximately $0.2$.
A.2 Zero-shot Question Accuracy after Finetuning
We evaluate performance using exact match accuracy, with results presented in Tables 3456, and 7. In the zero-shot setting, some tasks show accuracies significantly below chance rate (indicated in parentheses), as models occasionally generate responses outside the expected answer space. Notably, all models achieve zero-shot accuracies at or below chance rate across all tasks, suggesting that no superficial patterns were learned during fine-tuning to solve these tasks.
Appendix B Base Model Results
Figure 5 presents the attribution accuracies of the base models. While the overall attribution accuracies for linguistic pattern tasks are relatively low, the performance trends across different tasks and models exhibit similar patterns to those observed in the fine-tuned models (Figure 2).
Appendix C Prompt
We present sample of the exact prompt we used for our task, including the ones we used for testing attribution accurices and modified prompts for self-answer.
Normal Prompt
{mdframed}
[ backgroundcolor=black!5, topline=false, bottomline=false, rightline=false, leftline=false, innertopmargin=0.5em, innerleftmargin=0.5em, innerbottommargin=0.5em, innerrightmargin=0.5em, ]Input: they, Output: 6
Input: not, Output: 3
Input: I, Output: 5
Input: tell, Output: 7
Input: them, Output: 6
Input: were, Output: 6
Input: at, Output: 0
Input: yes, Output: 1
Input: right, Output: 9
Input: say, Output: 3
Input: they, Output:
Self-answer Prompt
{mdframed}
[ backgroundcolor=black!5, topline=false, bottomline=false, rightline=false, leftline=false, innertopmargin=0.5em, innerleftmargin=0.5em, innerbottommargin=0.5em, innerrightmargin=0.5em, ]<0>Input: they, Output: 6</0>
<1>Input: not, Output: 3</1>
<2>Input: I, Output: 5</2>
<3>Input: tell, Output: 7</3>
<4>Input: them, Output: 6</4>
<5>Input: were, Output: 6</5>
<6>Input: at, Output: 0</6>
<7>Input: yes, Output: 1</7>
<8>Input: right, Output: 9</8>
<9>Input: say, Output: 3</9>
<target>Input: they, Output: </target>
Among the 10 examples labeled <0> to <9>, select the single most helpful example for determining the answer to the <target> question. The correct answer to the target question is “6". To conclude this answer, we need to find one example that provides the necessary information. Therefore, the most helpful example is <
Appendix D Non-contrastive IA Methods Results
We further evaluate attribution accuracies using non-contrastive IA methods (specifically, non-contrastive GN, non-contrastive-I×$\times$×G, and non-contrastive IE), as shown in Figure 6. The results exhibit patterns that are highly consistent with those obtained from contrastive methods (Figure 2).
Appendix E More Results on Interpretability vs. Few-shot Number
Table 8910, and 11 show the Pearson correalation for Gemma-2-9B, Gemma-2-27B, Llama-2-7B and Llama-2-13B.
Zero-shot (%) Few-shot (%)
Associative-Recall $10.0$ $100.0$ ($10.0$)
Linear-or-Distinct $44.8$ $99.0$ ($50.0$)
Add-or-Multiply $11.8$ $100.0$ ($50.0$)
Verb-Object $0.0$ $100.0$ ($50.0$)
Tense-Article $0.0$ $100.0$ ($50.0$)
Pos-Title $0.0$ $98.0$ ($50.0$)
Table 3: The zero-shot and few-shot accuracy of the fine-tuned Gemma-2-2B model across all evaluation tasks. The chance rate is indicated in parentheses.
Zero-shot (%) Few-shot (%)
Associative-Recall $12.0$ $100.0$ ($10.0$)
Linear-or-Distinct $50.0$ $85.5$ ($50.0$)
Add-or-Multiply $12.0$ $100.0$ ($50.0$)
Verb-Object $0.0$ $94.8$ ($50.0$)
Tense-Article $0.0$ $100.0$ ($50.0$)
Pos-Title $50.0$ $98.8$ ($50.0$)
Table 4: The zero-shot and few-shot accuracy of the fine-tuned Gemma-2-9B model across all evaluation tasks. The chance rate is indicated in parentheses.
Zero-shot (%) Few-shot (%)
Associative-Recall $15.0$ $100.0$ ($10.0$)
Linear-or-Distinct $47.3$ $99.8$ ($50.0$)
Add-or-Multiply $48.5$ $97.3$ ($50.0$)
Verb-Object $0.0$ $99.5$ ($50.0$)
Tense-Article $50.0$ $100.0$ ($50.0$)
Pos-Title $45.5$ $97.8$ ($50.0$)
Table 5: The zero-shot and few-shot accuracy of the fine-tuned Gemma-2-27B model across all evaluation tasks. The chance rate is indicated in parentheses.
Zero-shot (%) Few-shot (%)
Associative-Recall $10.0$ $100.0$ ($10.0$)
Linear-or-Distinct $44.8$ $100.0$ ($50.0$)
Add-or-Multiply $11.8$ $99.0$ ($50.0$)
Verb-Object $0.0$ $100.0$ ($50.0$)
Tense-Article $0.0$ $100.0$ ($50.0$)
Pos-Title $0.0$ $98.0$ ($50.0$)
Table 6: The zero-shot and few-shot accuracy of the fine-tuned Llama-2-7B model across all evaluation tasks. The chance rate is indicated in parentheses.
Zero-shot (%) Few-shot (%)
Associative-Recall $10.0$ $100.0$ ($10.0$)
Linear-or-Distinct $50.0$ $99.8$ ($50.0$)
Add-or-Multiply $41.0$ $100.0$ ($50.0$)
Verb-Object $0.0$ $100.0$ ($50.0$)
Tense-Article $0.0$ $100.0$ ($50.0$)
Pos-Title $41.8$ $97.0$ ($50.0$)
Table 7: The zero-shot and few-shot accuracy of the fine-tuned Llama-2-13B model across all evaluation tasks. The chance rate is indicated in parentheses.
Task GN I×$\times$×G SA
Corr. $\Delta$ Corr. $\Delta$ Corr. $\Delta$
AR $-0.39$ $0.05$ $0.44$ $0.05$ $-0.93$ $0.16$
LD $1.00$ $0.35$ $0.89$ $0.07$ $-0.88$ $0.09$
AM $0.88$ $0.27$ $1.00$ $0.02$ $-0.89$ $0.08$
VO $0.99$ $0.13$ $-0.96$ $0.15$ $-0.88$ $0.10$
TA $-0.90$ $0.01$ $-0.73$ $0.11$ $-0.90$ $0.09$
PT $-1.00$ $0.07$ $-0.90$ $0.28$ $-0.93$ $0.09$
Table 8: TThe Pearson correlation coefficient between attribution accuracy and the number of few-shot examples for Gemma-2-2B.
Task GN I×$\times$×G SA
Corr. $\Delta$ Corr. $\Delta$ Corr. $\Delta$
AR $-0.07$ $0.07$ $1.00$ $0.41$ $-0.86$ $0.90$
LD $-0.75$ $0.07$ $0.00$ $0.04$ $-0.85$ $0.33$
AM $0.98$ $0.09$ $-0.93$ $0.02$ $-0.88$ $0.06$
VO $0.76$ $0.14$ $0.33$ $0.05$ $-0.85$ $0.14$
TA $0.86$ $0.36$ $-0.93$ $0.18$ $-0.90$ $0.09$
PT $0.90$ $0.22$ $0.65$ $0.06$ $-0.86$ $0.07$
Table 9: The Pearson correlation coefficient between attribution accuracy and the number of few-shot examples for Gemma-2-9B.
Task GN I×$\times$×G SA
Corr. $\Delta$ Corr. $\Delta$ Corr. $\Delta$
AR $0.28$ $0.06$ $-0.73$ $0.11$ $-0.86$ $0.29$
LD $0.97$ $0.24$ $-0.69$ $0.08$ $-0.89$ $0.09$
AM $0.98$ $0.26$ $-0.07$ $0.03$ $-0.85$ $0.37$
VO $-0.62$ $0.07$ $0.84$ $0.12$ $-0.86$ $0.15$
TA $0.96$ $0.19$ $-0.48$ $0.06$ $-0.92$ $0.10$
PT $0.90$ $0.14$ $0.25$ $0.04$ $-0.89$ $0.09$
Table 10: The Pearson correlation coefficient between attribution accuracy and the number of few-shot examples for Llama-2-7B.
Task GN I×$\times$×G SA
Corr. $\Delta$ Corr. $\Delta$ Corr. $\Delta$
AR $-0.74$ $0.06$ $0.32$ $0.07$ $-0.89$ $0.49$
LD $1.00$ $0.11$ $-0.85$ $0.07$ $-0.87$ $0.14$
AM $1.00$ $0.49$ $-0.97$ $0.14$ $-0.87$ $0.15$
VO $-0.81$ $0.04$ $-0.70$ $0.02$ $-0.84$ $0.43$
TA $0.99$ $0.13$ $0.98$ $0.04$ $-0.87$ $0.44$
PT $0.99$ $0.14$ $-0.87$ $0.02$ $-0.86$ $0.18$
Table 11: The Pearson correlation coefficient between attribution accuracy and the number of few-shot examples for Llama-2-13B.
Figure 5: Attribution accuracies for each task for base models.
Figure 6: Attribution accuracy using non-contrastive methods on fine-tuned models.